Find the total translational kinetic energy of
1.5 L of oxygen gas held at a temperature of
3◦C and a pressure of 2 atm.
Answer in units of J.
1
Expert's answer
2020-03-13T10:54:08-0400
"E=\frac{3}{2}nRT=\frac{3}{2}pV"

"E=\frac{3(2\cdot 101325)(1.5\cdot 10^{-3})}{2}=456\ J"